OpenStudy (ayee_ciera):

HELP!! MEDAL TO ALL THT HELP Two urns contain white balls and yellow balls. The first urn contains 2 white balls and 7 yellow balls and the second urn contains 3 white balls and 10 yellow balls. A ball is drawn at random from each urn. What is the probability that both balls are white?

OpenStudy (anonymous):

\[reqd.~probabilty=\frac{ 2 }{ 9 }*\frac{ 3 }{ 13 }=?\]

OpenStudy (ayee_ciera):

\[\frac{ 2 }{ 39 }\] thts wht i got

OpenStudy (ayee_ciera):

@surjithayer can you help me with some more????

OpenStudy (ayee_ciera):

can someone help me..?

OpenStudy (anonymous):

yw